OpenStudy (anonymous):

How many solutions are there to the following system of equations? 3x + 7y = –10 15x + 35y = –45

OpenStudy (anonymous):

answers could be 0 1 2 infinitely many

OpenStudy (zpupster):

the point the question is trying to make is parallel lines have no solutions and parallel lines have the same slope divide the 2nd eqaution by 5 what do you get

OpenStudy (anonymous):

i dont get it

OpenStudy (zpupster):

ok lets divide it

OpenStudy (anonymous):

ok

OpenStudy (zpupster):

15x/5+35y/5 = -45/5 can you do this??

OpenStudy (anonymous):

-9 right

OpenStudy (zpupster):

well 3x+ 7y = -9

OpenStudy (anonymous):

so there is no soultion?

OpenStudy (zpupster):

if yo subtract one equation from another you will end up with 0 or no solutions

OpenStudy (anonymous):

ok thanks!

OpenStudy (zpupster):

see no intersection --no solution

OpenStudy (zpupster):

you are welcome